OpenStudy (anonymous):

(x-y)(x²-2xy+y²) please help simplify

OpenStudy (anonymous):



OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@jdoe0001 @OrionsBelt

OpenStudy (anonymous):

vishnu can u help me

OpenStudy (anonymous):

you just distritube, and i already did that. i just wanna know if its right, and then i need to combine like terms.

OpenStudy (ranga):

Your second and fourth terms are off. The rest are correct.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

no wonder..

OpenStudy (anonymous):

can someone help me thru the rest of the problem

OpenStudy (anonymous):

so for the 2nd term, is it -2x^2y?

OpenStudy (ranga):

yes.

OpenStudy (anonymous):

and the 4th one?

OpenStudy (raden):

(x-y)(x^2 - 2xy + y^2) = x(x^2 - 2xy + y^2) - y(x^2 - 2xy + y^2) = x^3 - 2x^2y + xy^2 -x^2y + 2xy^2 - y^3 = x^3 - 3x^2y + 3xy^2 - y^3 alternative : remember that x^2 - 2xy + y^2 = (x-y)^2 so, (x-y)(x²-2xy+y²) = (x-y)(x-y)^2 = (x - y)^3 = x^3 - 3x^2y + 3xy^2 - y^3

OpenStudy (ranga):

fourth term is -x^2y

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

wait thats what i wrote..

OpenStudy (ranga):

exponent 2 is missing in the fourth term.

OpenStudy (anonymous):

oh nvm

OpenStudy (anonymous):

thats what i wrote on my paper, i just made a typo on the drawing up there^

OpenStudy (ranga):

ok.

OpenStudy (anonymous):



OpenStudy (anonymous):

yea i just forgot the 2. thanks

OpenStudy (ranga):

change the second term to -2x^2y too. Then combine the like terms.

OpenStudy (anonymous):

i did. thanks

OpenStudy (ranga):

you are welcome.